• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dalton

    Log entry time 18:34:22 on August 21, 2009

    Entry number 285160

    Followups:

    keyword=Day summmary


    After the hall was locked up at noon, a controlled access was made to
    work on the trigger.

    At about 15:00, the hall went to power permit to test the 9th dipole
    mapper. The mapper could be energised and moved through the magnet but
    consistently read zero both through the remote communication and on the
    face. A multiplexer was swapped out with one from Hall C as a test.
    This did not fix the problem and was backed out.

    At about 16:30, the hall was dropped into controlled access to reset LHRS
    quad 3 and RHRS quad 2 and 3. The first PSS interlock door at the hall
    entrance was worked on at this time too.

    The hall was again put into power permit and further work on the 9th
    dipole mapper was performed. The NMR electronics unit was swapped out
    with one from Hall C as a test. The new unit appears to work well in the
    shed, getting a lock on frequency and having a more promising scope
    trace, but the values in the software are still zeros. The next
    suggestion is to try replacing the MVME 712/M card. At this point the
    accelerator issues have taken precedence and work has stopped. The Hall
    C unit remains with us for the moment.



    A copy of this log entry has been emailed to: brads@jlab.org